# The Second Machine Turn: From Checking Proofs to Creating Concepts

### v1

@@
If we are to be active participants in the next revolution rather than mere bystanders, mathematicians must now do for the idea of “concept”; the units in which insights are measured, what Euclid once did for proof: give it a precise, mathematical formulation. Defining such a grammar - how concepts are discovered, compressed, ported, and retired - is not a task for engineers alone. It is a mathematical problem whose solution will determine whether future intelligences reason with us or past us. This essay takes a first step towards this goal. But to understand where we are headed, we must first understand where we have been. The story of mathematics has always been a story of co-evolution with our tools.
@@
Consider Euler’s famous claim that $1+2+3+\cdots=-1/12$. This wasn’t mysticism or error; it was an experimental observation. Euler had discovered that treating divergent series as if they converged led to consistent, useful results across multiple problems. The artifact was this surprising regularity. The explanation would come a century later with the theory of analytic continuation.
@@
The same is true in reverse. When mathematicians create new mathematics, they don’t proceed by systematic logical search. The reason is computational impossibility. Even a modest theorem might require a proof of twenty steps, each chosen from hundreds of possible moves. The search space explodes exponentially: a proof with merely $20$ lines could easily have at least $100^{20}$ possibilities, more than the number of atoms in the observable universe. No amount of computational power could explore such spaces by brute force.
@@
We follow the framework of [EJBL24] and view $Z$ as a corpus of structured data to be understood (current human mathematical output, formal proofs in ZFC, text on the internet, planetary orbits …) and $W$ as an internal, relational language modeling $Z$ through a structure preserving function $f$. We can then formalize the notion of compositional advantage through three critera:
First, the pattern must exhibit high expressive value: $K(Z)$ is large, meaning the mathematical corpus contains rich, complex phenomena worth explaining. The pattern of input-output relationships appeared in astronomy, physics, and pure mathematics, encoding deep regularities across diverse domains.
Second, the pattern must admit high reuse value: $K(Z|W)$ is small, meaning that the complexity of the data $Z$, given the conceptual library $W$ is small. Once ”function” was named, countless phenomena could be redescribed more concisely. The same abstract pattern (mapping inputs to outputs) compressed descriptions across all of mathematics.
Third, the pattern must possess semantic simplicity: $K(f)$ is small, meaning that the complexity of the mapping from $W$ to $Z$ is small. The rules for working with functions ( composition, inversion, restriction ) are remarkably simple yet generate endless complexity.
@@
The second approach pursues exploratory concept discovery, creating novel axiomatic frameworks and evaluating their generative power. Rather than extracting patterns from existing mathematics, the system would propose speculative frameworks $W^{\prime}$ and assess their mathematical value through metrics like unexpected connections: How many previously unrelated areas does $W^{\prime}$ unify?, theorem density: How many non-trivial results emerge from simple axioms? and pattern surprise: Using information theory, how much do discoveries in $W^{\prime}$ violate prior expectations?
This approach formalizes what Grothendieck achieved with schemes, creating an initially alien framework that proved its worth through the mathematics it enabled. The system would optimize
$L_{\text{explore}}=-\frac{K(Z_{\text{discovered}})}{K(Z_{\text{discovered}}|W^{\prime})}+\mu\cdot K(W^{\prime})$
@@
Recall our framework where mathematical knowledge consists of a relational language of concepts $W$ and their semantic interpretation $f:W\to Z$. The total complexity decomposes as:
$K(Z)=K(W)+K(f)+K(Z|W,f)$
Human mathematical practice reflects a severe constraint: our cognitive architecture drastically limits the complexity of $f$ we can execute. With working memory measured in single digits and computational speed in seconds per operation, we cannot mentally perform complex semantic mappings. We compensate by developing rich conceptual vocabularies (large $K(W)$) that enable us to reason with simple mental operations (small $K(f)$).
@@
The AI regime inverts this relationship. With $K(f)\gg K(W)$, these systems can maintain sparse conceptual frameworks while deploying semantic functions of astronomical complexity. Their understanding emerges through massive calculation and systematic search rather than conceptual insight. The proofs they generate may achieve absolute certainty while remaining opaque to human comprehension—correct at every step yet revealing no intelligible structure. The 10,000-page Riemann Hypothesis proof might use familiar number-theoretic concepts but connect them through computational paths no human could trace: each step simple, their composition incomprehensible. The AI hasn’t discovered alien mathematics but rather alien ways of doing familiar mathematics.

### v2

@@
If we are to be active participants in the next revolution rather than mere bystanders, mathematicians must now do for the idea of “concept”; the units in which insights are measured, what Turing did for computation: give it a precise, mathematical formulation. Defining such a grammar - how concepts are discovered, compressed, ported, and retired - is not a task for engineers alone. It is a mathematical problem whose solution will determine whether future intelligences reason with us or past us. This essay takes a first step towards this goal. But to understand where we are headed, we must first understand where we have been. The story of mathematics has always been a story of co-evolution with our tools.
@@
Consider Euler’s famous claim that 1+2+3+=1/121+2+3$\cdots$=-1/121 + 2 + 3 + ⋯ = - 1 / 12. This wasn’t mysticism or error; it was an experimental observation. Euler had discovered that treating divergent series as if they converged led to consistent, useful results across multiple problems. The artifact was this surprising regularity. The explanation would come a century later with the theory of analytic continuation.
@@
The same is true in reverse. When mathematicians create new mathematics, they don’t proceed by systematic logical search. The reason is computational impossibility. Even a modest theorem might require a proof of twenty steps, each chosen from hundreds of possible moves. The search space explodes exponentially: a proof with merely 202020 lines could easily have at least $100^{20}$ possibilities, more than the number of atoms in the observable universe. No amount of computational power could explore such spaces by brute force.
@@
We follow the framework of [elmoznino2024complexity] and view $Z$ as a corpus of structured data to be understood (current human mathematical output, formal proofs in ZFC, text on the internet, planetary orbits …) and $W$ as an explanation of $Z$ in an internal, relational language $p_{W}$ with the explanation determined by a structure preserving function $f:W\to Z$. We can then formalize the notion of compositional advantage through the following critera:
First, the data must exhibit high expressive value: $K(Z)$ is large, meaning the mathematical corpus contains rich, complex phenomena worth explaining. For example, the patterns of input-output relationships appearing in astronomy, physics, and pure mathematics encode deep regularities across diverse domains.
Second, the data must admit high reuse value: $K(Z|W)$ is small, meaning that the complexity of the data $Z$, given the conceptual library $W$ is small. Once ”function” was named, countless phenomena could be re-described more concisely. The same abstract pattern (mapping inputs to outputs) compressed descriptions across all of mathematics.
Third, the pattern must possess semantic simplicity: $K(f)$ is small, meaning that the complexity of the mapping from $W$ to $Z$ is small. Expressing planetary orbits in terms of functions is almost obvious. Conversely, when this correspondence grows more complicated as with modern algebraic geometry, mathematicians shifted to a new language of sheaves and schemes from varieties.
Fourth, the sentences $W$ must be typical in the distribution of $p_{W}$: $K(W|p_{W})$ is small, meaning that the sentences $W$ we care about are not especially cumbersome to express using $p_{W}$. The rules for working with functions (composition, inversion, restriction ) are remarkably simple in the language of set theory and nevertheless allow us to capture a very spectrum of ideas through the continuous, analytic and algebraic worlds amongst many others.
@@
The second approach pursues exploratory concept discovery, creating novel axiomatic frameworks and evaluating their generative power. Rather than extracting patterns from existing mathematics, the system would propose speculative frameworks $p_{W}$ and assess their mathematical value through metrics like unexpected connections: For how many previously unrelated areas $Z_{i}$ can we find explanations $W_{i}$ in the language $p_{W}$?, theorem density: How many non-trivial results emerge from simple axioms? and pattern surprise: Using information theory, how much do discoveries in $W$ violate prior expectations?
This approach attempts to formalize what Grothendieck achieved with schemes, creating an initially alien framework that proved its worth through the mathematics it enabled. The system would optimize
$L_{\text{explore}}=-\frac{K(Z_{\text{discovered}})}{K(Z_{\text{discovered}}|W)}+\mu\cdot K(p_{W})$
@@
Recall our framework where mathematical knowledge consists of a relational language of concepts $W$ in a language $p_{W}$ as well as a semantic interpretation $f:W\to Z$. When $W,p_{W}$ are chosen optimally and $Z$ is compositional, the total complexity decomposes as:
$K(Z)\sim K(p_{W},W,f)+K(Z|p_{W},W,f).$
We note in passing that the above approximate equality shows that $K(p_{W},W,f|Z)$ is logarithmically small in the size of $K(Z)$ and while at first site this might be surprising, it is explained by the fact that the optimal explanation of compositional data is indeed essentially determined by the data: the difficulty is in finding this explanation in practical terms.
As we have argued, finding such explanations is the primary goal of mathematical practice currently but we operate under a severe constraint: our cognitive architecture drastically limits the complexity of $f$ we can execute, as well as the complexity of $p_{W}$ that we can keep track of and internalize. With working memory measured in single digits and computational speed in seconds per operation, we cannot mentally perform complex semantic mappings and for the same reason, we also find it difficult to process languages that are contextually rich. We compensate by developing rich conceptual vocabularies (large $K(W)$) in modular languages (small $K(p_{W})$) that enables us to reason with simple mental operations (small $K(f)$).
@@
The AI regime inverts this relationship. With $K(f)\gg K(W)$, these systems can maintain sparse conceptual frameworks while deploying semantic functions of astronomical complexity. Their understanding emerges through massive calculation and systematic search rather than conceptual insight. The proofs they generate may achieve absolute certainty while remaining opaque to human comprehension; correct at every step yet revealing no intelligible structure. The 10,000-page Riemann Hypothesis proof might use familiar number-theoretic concepts but connect them through computational paths no human could trace: each step simple, their composition incomprehensible. The AI hasn’t discovered alien mathematics but rather alien ways of doing familiar mathematics.